OpenStudy (anonymous):

Find the value of the following expression: (38 ⋅ 2−5 ⋅ 90)−2 ⋅ 2 to the power of negative 2 over 3 to the power of 3, whole to the power of 4 ⋅ 328 (5 points) Write your answer in simplified form. Show all of your steps.

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

PEMDAS Parenthesis Exponents Multiplication/Division [left to right] Addition/Subtraction [left to right]

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

Is this your problem? \[[\left( 38.2-5.90 \right)-\left\{ \left( 2.2 \right)^{\frac{ 2 }{ 3 }} \right\}^3]^{4}.328\]

OpenStudy (anonymous):

are you here?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \left( 2.2 \right) ^{\frac{ 2 }{ 3 }*3}=\left( 2.2 \right)^2=4.84\]

OpenStudy (anonymous):

no its \[(3^{8} * 2^{-5} * 9^{0}) ^{-2} * (\frac{ 2^{-2} }{ 3^{-3} }) * 3^{28}\]

OpenStudy (anonymous):

i have to write it in simplified form

OpenStudy (anonymous):

\[=\left( \frac{ 3^8 }{ 2^5 }*1 \right)^{-2}*\frac{ 3^3 }{ 2^2 }*3^{28}\] \[=\frac{ 3^{8*-2} }{ 2^{5*-2} }*\frac{ 3^3*3^{28} }{ 2^2 }\] \[=3^{-16}\times 2^{10}\times3^3\times 3^{28}\times2^{-2}=3^{-16+3+28}\times2^{10-2}\] \[=3^{15}\times2^8\]

OpenStudy (anonymous):

this look so confusing @surjithayer